The owners of a candy store want to sell, for $6 per pound, a mixture of chocolate covered raisins, which usually sells for $3 per pound, and chocolate covered macadamia nuts, which usually sells for $8 per pound. They have a 60 pound barrel of the raisins. How many pounds of the nuts should they mix with the barrel of raisins so that they hit their target value of $6 per pound for the mixture?
Answer by stanbon(75887)

The owners of a candy store want to sell, for $6 per pound, a mixture of chocolate covered raisins, which usually sells for $3 per pound, and chocolate covered macadamia nuts, which usually sells for $8 per pound. They have a 60 pound barrel of the raisins. How many pounds of the nuts should they mix with the barrel of raisins so that they hit their target value of $6 per pound for the mixture?
-------------
Let amount of ccraisins be "x" lbs. ; value of those is 3x dollars
Amount of macnuts is "60-x" lbs. ; value of those is 8(60-x)=480-8x dollars
Amount of mixed nuts is 60 lbs. ; value of that is 6(60)=360 dollars
EQUATION:
3x+480-8x=360
-5x=-120
x=28 lbs (amount of chocolate covered raisins in the mixture)
60-x= 32 lbs ( amount of macadamia nuts in the mixture)
Cheers,
Stan H.